OpenStudy (anonymous):

log 6 (x + 4) + log 6 (x + 3) = 1

OpenStudy (anonymous):

log 6 (x^2+7x+12)=1 ==> x^2+7x+12=6 ===> x^2+7x+6=0 ===> (x+6)(x+1)=0 ===> x=-6 or x=-1